OpenStudy (anonymous):

HELP! As quick possible. Question : A hydrocarbon compound is burnt completely in air to form 7.2cm^3 of CO2 and 7.2 of water vapour. What is the molecular formula of the hydrocarbon compound? [ 24 dm^3 at room temp]

OpenStudy (anonymous):

Show working please

OpenStudy (callisto):

7.2 (unit?) of water vapour???

OpenStudy (anonymous):

7.2 cm^3

OpenStudy (callisto):

no. of mole of CO2 formed = (7.2/1000) / 24.0 = 0.0004 no. of mole of H2O formed = (7.2/1000)/24.0 = 0.0004 CxHy + (x+y/4)O2 -> xCO2 +(y/2) H2O x = 0.0004 y/2 = 0.004 => y = 0.0008 => CH2?

OpenStudy (callisto):

I think we can only get this but not the exact one, as there is insufficient information

OpenStudy (anonymous):

But the answer is C3H6

OpenStudy (callisto):

Is that all you've got for the question? No missing information like the molecular mass of the hydrocarbon?

OpenStudy (anonymous):

Nope. thats it. :(

OpenStudy (mos1635):

Callisto is right something is meesing with these gata all you can get is CxH2x

OpenStudy (anonymous):

My teacher says try solve it using the ratio. But i dont understand how.

OpenStudy (anonymous):

it has to be an alkene (CnH2n) coz the ratio of the products are balanced...the ratio of products divided by 2 gives the number of moles of hydrocarbon used...but u need the mass of the hydrocarbon to get the molecular formula?

OpenStudy (anonymous):

eg : if u take the complete combustion of C2H4 \[C_{2}H_{4} + O_{2} \rightarrow 2CO_{2} + 2H_{2}O\] number of moles of products equal = means the hydro carbon belongs to the "alkene" group Then u can find the no of moles of hydrocarbon used by divide the number of moles of Product divided by 2 = In your case 0.0003= no of moles of product formed, then u get the no of moles of hydrocarbon used by dividing t with 2 = u get 0.00015 Now the point is u need to know the Mass of hydrocarbon to get the molecular formula \[n= \frac{mass}{molecular formula}\]

OpenStudy (anonymous):

I got it now. Sorry because its an objective question. So, i guess we can pick according to the ratio. a. C2H5 b. C3H6 c. C3H8 d. C6H6 thanks everybody. sorry because i didnt include the answer :/ sorry

OpenStudy (anonymous):

its B:)

OpenStudy (mos1635):

yeap

OpenStudy (callisto):

only B is the compound that C:H is 1:2, as I've calculated before

OpenStudy (anonymous):

thts why it is B :P